OpenStudy (anonymous):

Find the product. (10^2)^3 is it 1000000?

OpenStudy (anonymous):

Yes it is..

OpenStudy (australopithecus):

(10^2)^3 = 10^6

OpenStudy (anonymous):

\[\large (x^a)^b = x^{a \times b}\]

OpenStudy (usukidoll):

yup. parenthesis first and then exponents.